OpenStudy (anonymous):

Emmett is graduating from college in six months, but he will need a loan in the amount of $2,350 for his last semester. He may either receive an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly, or his parents may get a PLUS Loan with an interest rate of 7.8%, compounded monthly. The Stafford Loan has a grace period of six months from the time of graduation. Which loan will have a higher balance and by how much at the time of repayment?

OpenStudy (tkhunny):

Silly problem. The Parent PLUS loan is due immediately upon disbursement. They have to start paying! It seems likely, then, that the PLUS Loan will be less at the time for graduation. Since we are not told what they will be paying, I think we're sunk.

OpenStudy (anonymous):

These are the choices i have. The PLUS Loan has a higher balance by $25.12. The Stafford Loan has a higher balance by $71.73. The PLUS Loan has a higher balance by $125.80. The Stafford Loan has a higher balance by $158.45.

OpenStudy (tkhunny):

Nice choices, but my objection stands. The naive, foolish, not-thinking-about-it-at-all answer might be: $2,350 for his last semester. He may either receive an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly, or his parents may get a PLUS Loan with an interest rate of 7.8%, c \(2350*(1 + 0.068/12)^{12} = 2514.88\) \(2350*(1 + 0.078/12)^{12} = 2540.00\) And \(2540.00 - 2514.88 = 25.12\)

OpenStudy (anonymous):

Thank you!

OpenStudy (tkhunny):

Hmmm... Still don't like the question.